Wite the slope intercept form of the equation of the line described
Through: (1,-4), parallel to y = -x + 3
Through: (2,1), perp. To y = 2x - 2

Found 2 solutions by solver91311, ewatrrr:
Answer by solver91311(24713)



Use these facts:






to determine the slopes of your desired lines, then use the point-slope form of an equation of a line:



where are the coordinates of the given point and is the calculated slope.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Line Through: (1,-4), parallel to y = -x + 3 ( m = -1) Slopes the same
New Line
y = -x + b ||Using pt(1,-4) to solve for b
-4 = -1 + b
-3 = b
y = -x - 3
Through: (2,1), perp. To y = 2x - 2 (m = 2) New Line, m = -1/2
New Line
y = (-1/2)x + b ||Using pt(2,1) to solve for b
1 = -1 + b
2 = b
y = (-1/2)x + 2